OpenStudy (anonymous):

Determine the value of alpha[0,2pi] for which sinalpha=cos(alpha/2)

OpenStudy (anonymous):

You have to convert sin(alpha) into cos.. Can you tell me how we can do this ??

OpenStudy (anonymous):

Do you know this : \[\sin(\alpha) = \cos(90^{\circ} - \alpha)\]

OpenStudy (anonymous):

Do you have keyboard with you @a11